### v1

# AutoPresent: Designing Structured Visuals from Scratch

Abstract
[…]
{strip}[Uncaptioned image]
Figure 1: Automatically generating slides from natural language instructions. We propose AutoPresent, a tool-augmented code generation method that follows natural language instructions to design slides from scratch, as shown in the examples. This allows for precise control over all elements, including textual content, images, visual layouts, coloring, and more.
1 Introduction
Designing structured visuals such as presentation slides from scratch is an essential skill for effective communication and conveying complex ideas [29]. Among various forms of visual communication, creating a compelling set of slides is a challenging problem, requiring content creation (text, pictures, diagrams, and more) and visual planning skills, to ensure the slides are well designed [24] and convey insights with clarity [3, 34]. Even human experts may need to spend hours iterating and polishing their slide decks [10] to produce high-quality designs with clear insights. While digital agents have demonstrated impressive capabilities in tasks such as software engineering [47], web navigation [52, 44], and free-form image design generation [8, 31], their creative capabilities in generating semi-structured communicative media like slide decks has not been extensively tested. Therefore, we ask: Can we employ powerful AI agents to create high-quality presentation slides that are well-structured and insight-revealing?
In this work, we formulate the natural language (NL) to slide generation task. At a high level, the user provides the system with a natural language instruction about the desired slide, and the system then generates an editable presentation, as shown in Figure 1. We consider three types of user instructions: (1) detailed instruction with images. (2) detailed instructions only. (3) high-level instructions, reflecting varying levels of design freedom.
[…]
To enable controlled and structured slide generation, we propose to create slides using program generation, where a model first generates a program from the natural language instruction, and then the program is executed to get the slide. We apply this approach to large language models (LLMs; LlaMa [11], GPT-4o [2]) and vision-language models (VLMs; LlaVa [49]). As illustrated in Figure 1, given a natural language instruction, the model first generates a Python program and then executes it to obtain a PPTX slide. We find that small models such as LlaMa (8B) and LlaVa (7B) are often unable to produce executable code. While GPT-4o can produce reasonable slides, it still exhibits a substantial gap in design quality compared to human-generated slides (§5). By further conducting iterative refinement, we find that models can self-refine and further improve slide quality. We also find that code generation approaches substantially outperform end-to-end image generation methods (Stable Diffusion [31], Dall-E [8]).
To further enhance the current model’s ability to generate high-quality slides, we present our open-sourced AutoPresent (8B) model (§4.2) which is fine-tuned from LlaMa 8B on the SlidesBench training set. AutoPresent achieves state-of-the-art performance among small open-sourced models and approaches the performance of the closed-sourced model GPT-4o. Since directly generating a long program is difficult for current models [14], we further create the SlidesLib library to simplify the program generation process. SlidesLib contains high-level functions that are basic such as add_title, and image-related such as search_image and generate_image. We show that LLMs and VLMs generally perform better when given access to SlidesLib.
Our main contributions can be summarized as follows:
[…]
  • We leverage NL-to-program generation methods with refinement to produce high-quality slides, and benchmark diffusion models, VLMs, and LLMs.
  • We train an 8B parameter open-source LLM, AutoPresent, that approaches the performance of GPT-4o, and design a programmatic tool library SlidesLib that facilitates slide program generation across models.
Figure 2: Illustration of SlidesBench. Each example of SlidesBench consists of three instructions: Detailed Instructions with Images, Detailed Instructions Only, and High-Level Instructions. The model is tasked to generate a slide based on the instruction, and the generated slide is evaluated on the metrics suite, which contains both the reference-free metrics and the reference-based metrics.
2 SlidesBench
In this section, we describe the creation of the SlidesBench benchmark. Each instance consists of a natural language instruction to create a slide, and the slide itself (in PPTX format) as a reference. SlidesBench includes three scenarios of varying difficulty levels designed to evaluate models with different user input. We describe the slide data collection (§2.1), three task setups (§2.2), and the annotation process (§2.3).
2.1 Slides Data Collection
We search the web and collect presentation slide decks from 10 domains, including art, marketing, environment, technology, etc. To select the highest-quality slide decks from each domain, we manually go through the relevant slide decks and conduct initial processing, by checking if all its slides (i) have visually structured layouts, and (ii) extractable media such as images (if any). For the slide decks with all slides satisfying (i) and (ii) in each domain, we incorporate one slide deck into the test set, and others into the training set. This results in a total of 10 and 300 slide decks (in PPTX format) for testing and training, each containing 20 slides on average. To respect the rights of the slide creators, we do not redistribute the slides. Instead, we provide a list of URLs for the slides that we used so that others can download the slides directly from the original website. We also provide an opt-out mechanism for any creator who does not want their slides in the dataset. We provide implementation details in §A.
2.2 Three Task Setups
We formulate the task as an NL-to-slide generation process. Given the reference slide, we curate three versions of natural language instructions, as shown in Figure 2, to represent slide generation tasks under varied difficulty levels. We introduce each setup below.
Detailed Instructions with Images  The first and easiest setting is to provide the models with all the necessary information and assets to produce the reference slide, including text and image content, formatting and layout specifications. This setting evaluates models’ visual planning abilities, such as arranging spatial layouts, maintaining formatting consistency, balancing content proportions, and emphasizing key elements.
Detailed Instructions Only  Since a user may not specify, or know exactly what images to put on a slide, we propose a detailed instruction only setting, where we provide the same natural language instruction provided in the detailed instruction with images setting, but replace the provided images with their natural language descriptions (e.g., “two people shaking hands”) generated by gpt-4o-mini. We then instruct the models to obtain the images using image searching or image generation tools. This setting further challenges models to interpret complex or compositional descriptions of images and obtain visuals that align with the slide context.
High-Level Instructions  In contrast to users who have a concrete target slide in mind and can spell out all detailed instructions, some users may only be able to express their needs on a high level. We thus devise a high-level instruction setting, where the natural language instructions are rather high-level and only provide a general topical idea of the slide, such as “create a title slide for Airbnb,” instead of detailing what logos and text to add and where, as exemplified in Figure 2. Models in this case need to both acquire or create content, and arrange the elements properly.
2.3 Example Annotation
To annotate the dataset, we collect natural language instructions paired with each slide. For each slide, we create three versions corresponding to the three setups in §2.2.
Detailed Instructions with Images  To produce detailed instruction with images, we use a scalable approach combining human-written examples and model-generated annotations. For each slide deck, we first write instructions for three example slides manually — including all necessary information (content, layout, formatting) to reproduce the slide, and providing paths to the images used in the slide (e.g., media/image_0.png), as shown in Figure 2 (top). We then use these (human-written instruction, reference slide) pairs as few-shot examples to prompt LLM (specifically, gpt-4o-mini) to generate natural language instructions for each slide in the current slide deck.***Including the three slides with human-written instructions, to ensure instructions for all slides are consistent in style and specificity. Further, for the test set, we manually examined and refined the instructions by correcting incorrect specifications, adding missing content, and removing unnecessary or untrue content.
Detailed Instructions Only  To produce detailed instruction only, we replace the image paths (e.g., media/image_0.png) with the natural language descriptions of the images(‘‘an artistic, colorful background’’). These descriptions are generated by gpt-4o-mini. For the test set, we manually refine the instructions to ensure that they do not refer to unavailable image paths (e.g., removing phrases like “use the provided images”), as shown in Figure 2 (middle).
High-Level Instructions  To create high-level instructions, we start with a similar approach by manually annotating three examples and then prompting the model to generate for all slides. Human-written instructions only provide a topical description of the slide and intentionally leave out specific content or layout details. This process ensures that the generated instructions remain concise and general, as shown in Figure 2 (bottom).
Overall, the instructions have an average of 115.6, 118.3, and 26.6 words under detailed instruction with images, detailed instruction only, and high-level instructions settings respectively, accompanied by an average of 1.1, 0.0, and 0.0 provided images.
3 Evaluation Metrics
In this section, we describe the evaluation metrics that we designed for SlidesBench. We propose two sets of evaluation metrics: reference-based metrics for measuring models’ instruction-following abilities (§3.1), and reference-free metrics to examine the design quality of model-generated content (§3.2). We also use executability to examine the success rate of each model (§3.3).
3.1 Reference-Based Metrics
Inspired by Design2Code [36] metrics, we implement four dimensions to examine the similarity between the model-produced slides and the reference slide.
Element matching  For the slide layout, we measure the total sizes of matched elements (in generated and reference slides) divided by the total sizes of all element, where each textbox, image, or shape constitutes an element. More concretely, we accurately parse out each element in the generated and reference slides, and compute their maximum matching using the match library.
Content similarity  For each pair of matched elements, we compute their content similarity. If the reference element is text, we calculate the textual similarity using the cosine similarity of the embeddings produced with sentence-transformer with the default all-MiniLM-L6-v2 model [28]. If the reference element is an image, we calculate the CLIP score [18] of the image in two elements. We report the average content similarity across all matched element pairs, if either element contains a non-empty text string or an image component.
Color similarity  We also measure the coloring similarity using the CIEDE2000 color difference formula [25], to quantify the perceptual difference between the colors. For every matched element pair, we measure the text font color similarity and element background color (if any). We additionally measure the color similarity between the background color of generated and reference slides.
Position similarity  In addition to content and formatting, we also calculate the positional similarity between each pair of matched elements. More concretely, we follow Si et al. [36] to normalize the element coordinates to $[0,1]$ by the slide page length and width. We compute the Manhattan distance between the elements and formulate positional similarity as $sim(r,g)=1-max(abs(x_{r}-x_{g},y_{r}-y_{g}))$.
Note that a low text, color, or position similarity score could come from differences in text, color, and positions, or derivative errors caused by the inaccurate element-matching process (e.g., it may match the title box in the generated slide to a content textbox in the reference slide, which has different content or coloring requirements).
3.2 Reference-Free Metrics
A well-designed slide generated by models may look very different from the reference slide. Therefore, we also propose four reference-free evaluation metrics, to independently assess the design quality of model-generated slides. To establish the metrics, we surveyed a wide range of literature on slide design principles [9, 32, 5, 42, 37], and summarized four major points as below and detailed in Table 1:
Metric Criteria
Text The title should be simple and clear to indicate the main point. For main content, avoid too many texts and keep words concise. Use a consistent and readable font size, style, and color.
Image Use high-quality images with a reasonable proportion.
Layout Elements should be aligned, do not overlap, and have sufficient margins to each other. All elements should not exceed the page.
Color Use high-contrast color especially between the text and the background. Avoid using high-glaring colors.
Table 1: Reference-free metrics, all evaluated in 0-5 scale.
Text  Using concise texts is important for slides to engage with the audience. An ideal slide should have a clear title, concise main content, and readable formatting.
Image  Using appropriate visuals can engage audiences. We hence measure if models can find high-quality images and properly use them to enhance the slide quality.
Layout  Slide layout is crucial to create visual balance. We examine whether all elements are within the slide, have no overlap, and align properly with the relevant elements.
Color  Vivid and consistent color use in slides can help deliver insights. We check if the slide uses high-contrast colors to facilitate visibility, and avoid high-glaring colors to discourage user engagement.
Validation of Reference-Free Evaluation
For all the metrics, we provide the image version of the slide and ask the gpt-4o model to produce a score between 0–5. To examine the reliability of this model-based evaluation, we conduct a human study and compare the intraclass correlation coefficient (ICC) between two human annotators and model evaluation, on all ground-truth slides. Our examination gives high ICC scores across all four metrics: $73.8\%$$85.3\%$, which are well within the range of what is typically considered “high agreement”. In experiments in later sections, we scale these 0-5 scale scores to the 0–100 range to enable comparisons on this more standard scale. We still evaluate with 0-5 scale to maintain a robust, human-aligned evaluation process.
3.3 Executability
Particularly for methods based on code generation (§4.1), we additionally measure the execution success rate to account for invalid programs. Concretely, we count the percentage of successfully executing programs generated by models among all examples. We report reference-based and free scores for executable slides only, to fairly compare their design quality. But we report ‘Overall’ scores for all slides by assigning zeros to non-executing slides, to account for execution failures. We report all metrics for successfully executing and all slides in §E.
4 Method
[…]
4.1 Slides via NL-to-Code Generation
Generating Python Programs
Given natural language instructions in §2, models are tasked with generating Python programs using publicly available libraries such as python-pptx. The model receives two (natural language instruction, Python program) pairs as in-context examples, followed by the test instruction, and generates a Python program which is then executed and will ideally yield a PPTX file containing the requested slide.
Generating Programs with SlidesLib
Nonetheless, the programs above could be very long and complex (170 lines on average), which could be challenging for models to generate entirely correctly, as shown in previous work [14]. To address this, we design SlidesLib, a library that provides easier-to-use interfaces for several common actions such as setting a title or setting background color. Using SlidesLib, the average program length is reduced to 13 lines, significantly easing the generation task. As shown in Table 2, SlidesLib includes 4 functions for basic operations and 3 functions for image search and generation, these functions allow models to produce more concise and modular programs. To enable the model to generate programs using SlidesLib, we follow the visual programming method [41] by providing a prompt that includes the documentation of the functions and two in-context examples. See more SlidesLib details in §B.
Function Description
add_title Insert a title in the slide.
add_text Insert text at a specific location.
add_bullet_points Insert a textbox with bullet points.
add_image Insert image at a specific location.
generate_image Call an image generator (Dall-E 3) given a query.
search_image Search for an image on a search engine (Bing).
search_screenshot Display a query on a web browser (Google Chrome) and take a snapshot of the search result.
Table 2: Basic (top) and image-specific (bottom) functions provided by SlidesLib.
Method ExecutionReference-Based Reference-Free Overall
element content color position text image layout color
Reference 100.0 59.7 81.5 73.5 65.7
End-to-end Image Generation
Stable-Diffusion* 100.0 74.5 33.4 9.0 75.0 19.6 45.1 36.9 40.5 48.0
DALLE 3* 100.0 75.5 39.9 9.2 76.1 32.7 87.3 56.7 53.4 50.2
Code Generation w/o SlidesLib
LLaVA (7B) 11.3 61.9 97.3   6.2 70.8 41.6 100.0 29.2 25.7   6.1
LLaMA (8B)   2.1 74.0 94.6 12.5 81.2 50.0   8.3 50.0 50.0   1.3
GPT-4o 89.2 83.3 91.6 10.5 77.0 51.9 72.8 53.7 54.7 55.1
AutoPresent (ours) 79.0 67.7 79.7 10.9 75.9 45.3   62.7 54.2 60.9 45.2
Code Generation w/ SlidesLib
LLaVA (7B) 20.0 80.5 80.5   3.5 64.0 37.5 48.0 29.5 43.5   9.7
LLaMA (8B) 54.4 78.3 91.2 7.5 69.5 46.0 68.2 47.6 53.1 33.5
GPT-4o 86.7 86.2 92.5 12.7 76.3 54.6 83.7 70.5 59.4 58.0
AutoPresent (ours) 84.1 84.2 92.2 18.1 67.2 47.8 73.2 58.6 64.7 55.0
Table 3: Results with detailed instructions with images. We found that small models like LlaVa (7B) and LlaMa (8B) can barely generate any slides, while AutoPresent (8B) generates slides on par with GPT-4o. All the models still underperform humans.
4.2 AutoPresent
Using the slides in the training set of SlidesBench, we construct (natural language instruction, program) pairs to form training data to train an open-sourced 8B model, AutoPresent. This model is based on the LlaMa-3.1-8B-Instruct and trained using LoRA [20] with a rank of 128.
Training Data Construction
To create (natural language instruction, program) training pairs, we generate two versions of canonical program solutions for each slide:
(i) Basic Python Programs  We derive canonical programs (that is, programs that can be executed to reproduce the slide) without SlidesLib. To do this, we manually design an extraction script that (i) extracts each element (e.g., text and image) on the given slide, and (ii) produces a rule-based program that adds each element to the slide. After extracting and adding each element to the slide, the resulting program accurately reproduces the original slide.
(ii) SlidesLib Python Programs  We also generate canonical programs using SlidesLib, by transforming snippets from the programs above into SlidesLib function calls. To reproduce images in detailed instruction only and high-level instructions settings, we generate a short caption for each image and provide it to GPT-4o to generate the program for producing that image using search_image or generate_image functions. More details of this automatic program generation process are in §B.2.
Training Set Composition
[…]
  1. 1.
    (detailed instruction with images, python program)
  2. 2.
    (detailed instruction with images, SlidesLib program)
  3. 3.
    (detailed instruction only, SlidesLib program)
  4. 4.
    (high-level instructions, SlidesLib program)
These training sets allowed us to train four specialized models that address different challenges, which we report in Table 3 (1,2) and Table 4 (3,4).
4.3 Iterative Refinement
Slide generation is by nature an iterative process and often requires visual-based refinements after the first draft. To enable models to refine slides as humans do, we explore an iterative refinement procedure, where the model is tasked to self-refine the slide it generated. Specifically, in the setting using SlidesLib, we provide GPT-4o (capable of consuming slide images) with the original language instruction, the program it generated in the first pass, and a snapshot of the rendered slide; the model is then asked to generate a new program based on these information to refine the slide quality by tweaking colors, spacing, and other aspects of the slide. See the prompts of this process in §D.
5 Experiments and Results
Figure 3: Examples of slides generated by different methods in three scenarios. End-to-end image generation methods fail to generate structured and clear slides. Small open-sourced models like LlaMa and LlaVa can barely generate any usable slides, while AutoPresent produces quality slides. Adding SlidesLib improves GPT-4o’s performance on detailed instruction only and high-level instruction tasks.
We first introduce the experimental setup (§5.1), then present the results under various scenarios (§5.2).
5.1 Experimental Setup
Code Generation Approaches
For code generation approaches, we sample $n=3$ responses and iteratively go through them, using the first successfully executing program as the final output of the model. If none of the $n$ responses execute successfully, we count it as an execution failure. In addition to AutoPresent4.2), we benchmark several LMs out-of-the-box, including open-weights Llama 3.1 (8B, Instruct), the vision-language LlaVa v1.5 model (with a Vicuna-7B-v1.5 LM backbone); and the proprietary GPT-4o model (the gpt-4o-2024-08-06 checkpoint).
End-to-End Image Generation
We compare code generation with end-to-end neural image generation methods, which are a common way to produce visuals. These methods are good at creating scenic or artistic images, but may be imprecise in content (esp. text) and do not support easy further modification by users. We benchmark Stable-Diffusion 2 [31] and DALL-E 3 [8] by asking them to output slides given the natural language instructions. We adjust our reference-based evaluation procedure by first segmenting slide images into elements using Tesseract OCR[38] and further parse out the texts of the elements, then applying the default calculation process as in §3. For the detailed instruction with images setting, we also report the results of the end-to-end image generation methods, marked with a “*” to indicate that they do not actually use the image inputs.
Method Detailed Instructions Only High-Level Instructions
exec ref-based ref-free overall exec ref-based ref-free overall
End-to-End Image Generation
SD2 100.0 48.0 35.5 48.0 100.0 47.7 31.5 47.7
DALLE 3 100.0 50.2 57.5 50.2 100.0 50.7 53.6 52.2
Code Generation w/o Library
LLaVA 17.9 56.9 47.4   9.3 19.5 50.2 47.3   9.5
LLaMA   4.6 61.4 35.1   2.8   8.7 55.6 50.1   4.8
GPT-4o 50.3 66.8 50.0 28.7 70.8 60.3 57.0 39.7
Code Generation w/ Expert-Designed Library
LLaVA 17.4 58.2 33.8   8.0 25.1 50.1 36.7 10.9
LLaMA 60.5 61.7 56.6 37.4 76.9 56.8 58.3 43.7
GPT-4o 87.7 64.2 65.8 56.3 97.4 60.1 71.2 58.5
AutoPresent 89.2 61.9 58.7 55.2 86.6 55.2 61.5 47.8
Table 4: Results under detailed instruction only and high-level instructions settings. We assign 100% execution success rates for all end-to-end image generation methods because they do not generate programs and would not suffer from execution errors.
5.2 Quantitative Results and Analysis
Table 3 shows the result of detailed instruction with images scenario and Table 4 shows the result of detailed instruction only and high-level instructions scenarios.
In the top row of Table 3, we first measure the scores of the reference slides, which shows that the quality of the human-created slides is among the highest.
Compared to the scores achieved by GPT-4o, smaller open-source models such as Llama 3.1 and LlaVa barely produce any working slides out of the box. Although the significant gaps of 49.949.949.949.955.055.055.055.0 points exist in the detailed instruction with images setting, this gap shrinks to 22.222.222.222.234.634.634.634.6 when no visuals are provided a priori, in detailed instruction only and high-level instructions scenarios (Table 4). This demonstrates significant challenges in obtaining images in slides. In contrast to the low performance of open-weight models out-of-the-box, AutoPresent’s performance approaches that of GPT-4o.
End-to-End Image Generation When no visuals are provided, end-to-end image-generation methods perform worse than the best code-generation approaches in both the reference-based and reference-free metrics, especially in generating accurate content. These methods also often produce creative figures without aligning with the design principles of slides, indicating its poorer controllability.
Effect of LibrarySlidesLib brings observable gains in LlaMa and LlaVa in all three scenarios by at most $34.0$ points; and similarly increases the strong GPT-4o performance across scenarios, especially when no images are provided. This suggests the benefits of generating more modular and concise programs for structured visual design.
VLM vs. LLM  When no helper functions are presented, the one VLM that we tested (LlaVa) outperforms its LLM counterpart LlaMa in all scenarios by 5.15.15.15.17.57.57.57.5 points. However, LlaVa shows limited ability in using functions presented in context, as demonstrated by the large margin the library-augmented LlaMa has over LlaVa (12.112.112.112.126.226.226.226.2). All LLMs (LlaMa, GPT-4o) perform worse when the instructions become less specified (detailed instruction with images $\rightarrow$ detailed instruction only $\rightarrow$ high-level instructions). Nonetheless, SlidesLib can greatly mitigate this degradation due to the loss of input specificity, and help models produce better outcomes across all three scenarios.
5.3 Qualitative Case Study
We illustrate several models-produced slides in Figure 3. For end-to-end image generation methods, the design is more creative and often more attractive, but the text does not constitute meaningful words, or even the characters themselves are not valid.
On the other hand, code generation methods, especially weaker LlaMa and LlaVa models, suffer more from visual layout — elements often overlap with each other or exceed the canvas, making it challenging for the audience to obtain all information clearly.
In contrast, AutoPresent generates slides with appropriate layouts without undesirable element overlaps. In addition, they better follow the user instructions and are not overly creative like the image generation methods.
5.4 Perceptual Evaluation
Model Pairs Detailed+Images Detailed Only
t-stat p-val t-stat p-val
(GPT-4o, LlaMa) 13.206 0.000 8.630 0.000
(AutoPresent, LlaMa) 13.180 0.000 2.955 0.004
(GPT-4o, AutoPresent) -0.445 0.657 8.203 0.000
Table 5: Paired t-test results comparing model performance across detailed instruction only setting and detailed instruction with images setting. AutoPresent and GPT-4o outperforms LlaMa with a statistically significant difference in both settings.
We performed a qualitative evaluation on 10 randomly selected slides from each domain generated by GPT-4o, Llama-8B, and AutoPresent under the detailed instruction with images and detailed instruction only settings. We also add the ground-truth reference slide to evaluate the performance gap between current models and human slide creators. We shuffle these slides and ask the annotators to rank each slide from 1-5 based on how likely they would be to use the slide. For the detailed instruction with images setting, we collect 25 responses in total, and for detailed instruction only, we collect 16 responses in total. We provide more details of the evaluation process in §F.
The result is shown in Figure 4. By performing the paired $t$-test, we found differences between the models pairs in terms of user preferences, as shown in Table 5: (1) In both settings, AutoPresent and GPT-4o perform statistically significantly better than LlaMa. (2) In detailed instruction with images setting, GPT-4o and AutoPresent has no significant differences (3) In the detailed instruction only setting, AutoPresent is slightly worse than gpt-4o, aligning with our quantitative evaluations in Table 4. All three models still have an overall performance gap compared with human-designed slides, indicating room for improvement on the slide generation task.
Figure 4: Perceptual evaluation results on detailed instruction (1) with images and (2) only settings. We ask the users to score the quality of each slide from 1-5 and report the average score of each model. The user reported preference on GPT-4o and AutoPresent compared with LlaMa, while still having a gap with human-designed slides.
5.5 Result after Iterative Refinement
Method Detailed + Images Detailed Only High-Level
GPT-4o 58.0 56.3 58.5
Refinement 59.5 59.5 59.8
Table 6: Overall scores after applying refinement in the three scenarios, demonstrating that refinement boosts performance in all three scenarios, especially the detailed instructions only task.
Finally, as shown by Table 6, we find that refinement improves model performance on all three challenges, especially in the detailed instruction only setup by $3.2$ points. We present representative cases in Figure 5, which indicates that refinement can improve content layout and detailed controls on coloring and sizing.
Figure 5: Auto-refinement results with GPT-4o, where the model further addresses some previously neglected instructions (marked in green), such as shape, background color, and text.
6 Related Work
Language and Vision Model-Based Agents  Agents based on large language models (LLMs) [2, 11] and vision-language models (VLMs) [4, 49] have been widely adopted in various tasks such as web navigation [50, 52, 23], software engineering [47, 48], and web development [26, 36]. Creation of presentation materials is another common task [10] that has both similarities and differences from these more widely examined tasks.
Generating Programs for Vision Tasks  End-to-end image generation models such as diffusion [19, 31, 46] and GAN [15, 33] are widely used at producing scenic images, yet falling short on more structured visuals such as websites and slides [36]. Generating programs (i.e., image-editing actions) is a useful means to get structured visuals [17, 41, 39, 43, 14], including Tikz figures [6, 7], SVG [30, 35], posters [49], and user interfaces [36, 27]. However, they often require detailed inputs and are limited to specific, simple figure types, so they are still far from creating complex, editable presentation slides from scratch. Our work extends this line of research by formulating and benchmarking the natural-language-to-slide generation task.
Automatic Slide Generation  Previous works on slide creation mostly focus on basic extraction from provided documents [22, 21, 34, 40, 12] or having models generate content given a topic [3, 45, 1] without addressing how to organize content visually. More recently, some benchmarks [16, 51] and methods [13] have emerged that follow detailed instructions for slide editing (e.g., adjust the font size of the title from 20 to 24) of an existing slide. In contrast, we synthesize more complex and structured programs that can generate slides from scratch, including content creation, visual arrangement, and fine-grained editing, instead of refining an existing slide.
7 Conclusion and Limitations
In this work, we address the challenge of creating structured visuals from scratch. Specifically, we introduced SlidesBench, the first benchmark for automatic slide generation with evaluation metrics based on and free of reference slides. We benchmark multiple end-to-end image and program generation approaches, and demonstrate that AutoPresent with SlidesLib achieves comparable performance with the top GPT-4o model. Our further exploration in iterative refinement also reveals certain effectiveness in self-refinement. This work is an initial step towards automated generation of structured visuals. Specifically, it focuses on single-slide generation and produces full slide code in a single pass, without leveraging iterative design workflows. Future research could address these limitations by expanding to full slide decks, adopting gradual and interactive slide generation, and incorporating slide-specific features like animations. Further, integrating more design principles, such as optimizing for attention capture and information clarity, would be crucial for making generated slides more impactful and effective.
Acknowledgment
We would like to thank Yutong Bai for helping us draft Figure 1 and providing feedback on the paper, David Chan for providing detailed suggestions on the introduction, and Frank Xu and Sean Welleck for discussions at the initial stage of this project. Junyi Zhang and Haven Feng for feedback on the project.
Appendix A SlidesBench Details
A.1 Slide Deck Domains
The 10 domains we cover in SlidesBench include:
  1. 1.
    Art Photos
  2. 2.
    Business
  3. 3.
    Career
  4. 4.
    Design
  5. 5.
    Entrepreneur
  6. 6.
    Environment
  7. 7.
    Food
  8. 8.
    Marketing
  9. 9.
    Social Media
  10. 10.
    Technology
A.2 Slide Deck Source
There existis large amount of slide decks on the internet including Google Search, Bing Search etc. For convenience, we collect a list of slides from the slideshare.com website.
Appendix B SlidesLib Details
In this section, we provide the detailed documentation and examples for all functions in our SlidesLib.
B.1 SlidesLib Implementation
Figure 6 shows the basic functions and Figure 7 shows the image-oriented functions.
add_title(slide, text, font_size, font_color, background_color)
"""Addatitletexttotheslidewithcustomfontsizeandfontcolor(RGBtuple).
Args:
␣␣␣␣slide:Slideobjectasinpptxlibrary
␣␣␣␣text:str,Titletexttobeadded
␣␣␣␣font_size:int,Fontsizeinint(pointsize),e.g.,44
␣␣␣␣font_color:tuple(int,int,int),RGBcolor,e.g.,(0,0,0)
␣␣␣␣background_color:Optional,tuple(int,int,int),RGBcolor,e.g.,(255,255,255)
Rets:
␣␣␣␣slide:Slideobjectwiththetitleadded
add_text(slide, text, coords, font_size, bold, color, background_color, auto_size)
"""Addatextboxataspecifiedlocationwithcustomtextandcolorsettings.
Args:
␣␣␣␣slide:Slideobjectasinpptxlibrary
␣␣␣␣text:str,Texttobeadded
␣␣␣␣coords:list(float),[left,top,width,height]ininches
␣␣␣␣font_size:int,Fontsizeinint(pointsize),e.g.,20
␣␣␣␣bold:bool,Trueifbold-typethetext,Falseotherwise
␣␣␣␣color:tuple(int,int,int),RGBcolor,e.g.,(0,0,0)
␣␣␣␣background_color:Optional,tuple(int,int,int),RGBcolor,e.g.,(255,255,255)
␣␣␣␣auto_size:bool,Trueifauto-sizethetextbox,Falseotherwise
Rets:
␣␣␣␣slide:Slideobjectwiththetextboxadded
add_bullet_points(slide, bullet_points, coords, font_size, color, background_color)
"""Addatextboxwithbulletpoints.
Args:
␣␣␣␣slide:Slideobjectasinpptxlibrary
␣␣␣␣bullet_points:list(str),Listoftextstobeaddedasbulletpoints
␣␣␣␣coords:list(float),[left,top,width,height]ininches
␣␣␣␣font_size:int,Fontsizeinint(pointsize),e.g.,18
␣␣␣␣color:tuple(int,int,int),RGBcolor,e.g.,(0,0,0)
␣␣␣␣background_color:Optional,tuple(int,int,int),RGBcolor,e.g.,(255,255,255)
Rets:
␣␣␣␣slide:Slideobjectwiththebulletpointsadded
add_image(slide, image_path, coords)
"""Addanimageintheprovidedpathtothespecifiedcoordsandsizes.
Args:
␣␣␣␣slide:Slideobjectasinpptxlibrary
␣␣␣␣image_path:str,Pathtotheimagefile
␣␣␣␣coords:list(float),[left,top,width,height]ininches
Rets:
␣␣␣␣slide:Slideobjectwiththeimageadded
set_background_color(slide, color)
"""Setbackgroundcolorforthecurrentslide.
Args:
␣␣␣␣slide:Slideobjectasinpptxlibrary
␣␣␣␣color:tuple(int,int,int),RGBcolor,e.g.,(255,255,255)
Rets:
␣␣␣␣modifiedslideobject
Figure 6: Documentation for the basic functions in our SlidesLib.
google_search_screenshot(question, save_path)
"""SearchaquestiononGoogle,andtakeascreenshotofthesearchresult.
Savethescreenshottosave_path,andreturnthepath.
Args:
␣␣␣␣question:str,ThequestiontosearchonGoogle.
␣␣␣␣save_path:str,Thepathtosavethescreenshot.
Returns:
␣␣␣␣Thepathofthesavedscreenshot.
search_image(query, save_path)
"""SearchforanimageonGoogleanddownloadtheresulttosave_path.
Args:
␣␣␣␣query:str,Thequerytosearchfor.
␣␣␣␣save_path:str,Thepathtosavethedownloadedimage.
Rets:
␣␣␣␣thesave_path.
generate_image(query, save_path)
"""Generateanimageusingdiffusionmodelbasedonatextquery,andsavetheimagetothepath.
Args:
␣␣␣␣query:str,Thetextquerytogeneratetheimage.
␣␣␣␣save_path:str,Thepathtosavethegeneratedimage.
Rets:
␣␣␣␣Thepathofthesavedimage
Figure 7: Documentation for the image-oriented functions in our SlidesLib.
B.2 SlidesLib Usage Example
Figure 8 shows two example programs using multiple SlidesLib functions to produce slides.
# Create slide with the title ’NLP Can Answer Questions’ in large, bolded font in the top center of the page. Below it, put a screenshot of the google search result of the question ’Where was the first movie theater in the U.S?’ in the middle of the page.
from pptx import Presentation
from pptx.util import Inches, Pt
from library import add_text, google_search_screenshot, add_image
presentation = Presentation()
presentation.slide_width = Inches(16)
presentation.slide_height = Inches(9)
slide_layout = presentation.slide_layouts[0] # choose a layout template
slide = presentation.slides.add_slide(slide_layout)
add_text(slide, "NLPCanAnswerQuestions", coords=(1, 0.5, 8, 1), font_size=36)
img_path = google_search_screenshot("WherewasthefirstmovietheaterintheU.S?", save_path="screenshot.png")
add_image(slide, "screenshot.png", coords=(2.5, 2, 6, 4))
presentation.save("target_path.pptx")
# Create a slide titled ’Interior Design’ in bold, dark-green color in the center of the page. For the background, consider using a picture with a color, artistic vibe, ensure enough contrast between the colors of text and background.
from pptx import Presentation
from pptx.util import Inches, Pt
from library import generate_image, add_image, add_text
presentation = Presentation()
presentation.slide_width = Inches(16)
presentation.slide_height = Inches(9)
slide_layout = presentation.slide_layouts[5] # choose a layout template
slide = presentation.slides.add_slide(slide_layout)
background_img = generate_image("Ancolorful,artisticbackground", "colorful.png")
add_image(slide, "colorful.png", coords=(0.0, 0.0, 16, 9))
add_text(slide, ’InteriorDesign’, coords=(0.0, 2.4, 13.3, 1.3), font_size=80, bold=True, color=(0, 0, 0), background_color=(255, 255, 255), auto_size=True)
presentation.save("path.pptx")
Figure 8: Example programs to produce slides using SlidesLib.
Appendix C Training Details for AutoPresent
The training parameters for AutoPresent are summarized in Table 7.
Parameter Value
LoRA Parameters
LoRA rank 128
LoRA alpha 32
LoRA dropout 0
Random state 3407
RS-LoRA Disabled
LoFT-Q config None
Trainer Parameters
Batch size (per device) 1
Gradient accumulation steps 2
Warmup steps 20
Epochs 1
Learning rate 3e-4
Mixed precision FP16
Weight decay 0.01
Scheduler Linear
Seed 3407
Table 7: Training details for AutoPresent. LoRA and Trainer parameters are described in detail.
Appendix D Refinement Details
We provide the prompts that we used for auto-refinement in Figure 9. We input the instruction and the in-context examples, the previous code generated by the model, and the snapshot of the slide generated by executing this code to the model and let it correct the code.
Youareanexpertpresentationslidesdesignerwhocreatesmodern,fashionable,andstylishslidesusingPythoncode.YourjobistogeneratetherequiredPPTXslidebywritingandexecutingaPythonscript.Makesuretofollowtheguidelinesbelowanddonotskipanyofthem:
1.␣␣Ensureyourcodecansuccessfullyexecute.Ifneeded,youcanalsowriteteststoverifyyourcode.
2.Maintainproperspacingandarrangementsofelementsintheslide:makesuretokeepsufficientspacingbetweendifferentelements;donotmakeelementsoverlaporoverflowtotheslidepage.
3.Carefullyselectthecolorsoftext,shapes,andbackgrounds,toensureallcontentsarereadable.
4.Theslidesshouldnotlookemptyorincomplete.Whenfillingthecontentintheslides,maintaingooddesignandlayout.
Followtheinstructionbelowtocreatetheslide.
Iftheinstructionislongandspecific,followtheinstructioncarefullyandaddallelementsasrequired;
ifitisshortandconcise,youwillneedtocreatesomecontent(text,image,layout)andimplementitintotheslide.
Ifyouneedtousetheprovidedimages,refertotheimagefilenamesintheinstructions.
Finally,yourcodeshouldsavethepptxfiletopath"output.pptx"
APILibraries:
#INSERT_API_DESCRIPTIONS_HERE
##Examples
#INSERT_IN_CONTEXT_EXAMPLES_HERE
ModificationTask:
Instruction:INSERT_INSTRUCTION_HERE
PreviousCode:
INSERT_PREV_CODE_HERE
SlideSnapshot:Seeimage.
Task:Basedontheobserveddrawbacksintheprovidedslideimage,modifytheexistingcodeaccordinglytoimprovetheslide’sdesignandfunctionality.
Yourmodification:
defgenerate_presentation():
Figure 9: Prompt we used for Auto-Refinement. The model receives the APIs and instruction, the previous generated slide and code, and is tasked to re-write the code to do slide refinement.
Appendix E Detailed Results
We report two sets of evaluation metrics (reference-based and reference-free) in both their average value on all slides (i.e., un-weighted by execution success) and on successfully rendered slides (i.e., weighted by execution success).
E.1 Detailed Instructions with Images
Table 8 shows all metrics down-weighted by the execution success rate; Table 9 shows reference-based and reference-free metrics without down-weighting by execution success.
Method ExecutionReference-Based Reference-Free Average
block text color position text image layout color
Human 100.0 - 59.7 81.5 73.5 65.7 -
Code Generation w/o Library
LLaVA (7B) 11.3   7.0 11.0   0.7   8.0   4.7 11.3   3.3   2.9   6.1
LLaMA (8B)   2.1   1.5 1.9   0.3   1.7   1.0   0.2   1.0   1.0   1.3
GPT-4o 89.2 74.3 80.7   9.4 68.7 46.3 64.9 47.9 48.8 55.1
AutoPresent (ours) 79.0 53.5 63.0   8.6 60.0 35.8 49.5 42.8 48.1 46.3
Code Generation w/ Expert-Designed Library
LLaVA (7B) 20.0 16.1 16.1   0.7 12.8   7.5   9.6   5.9   8.7   9.7
LLaMA (8B) 54.4 42.6 49.6   4.1 37.8 25.0 37.1 25.9 28.9 33.5
GPT-4o 86.7 74.7 80.2 11.0 66.1 47.3 72.5 61.1 51.4 58.0
AutoPresent (ours) 84.1 70.8 77.5 15.2 56.5 40.2 61.6 49.3 54.4 55.0
Table 8: Slide generation results (weighted by execution success) under the detailed instructions with images scenario.
Method ExecutionReference-Based Reference-Free Avg
block text color pos text img layout color
Human 100.0 - 59.7 81.5 73.5 65.7 -
Code Generation w/o Library
LLaVA (7B) 11.3 61.9 97.3   6.2 70.8 41.6 100.0 29.2 25.7   6.1
LLaMA (8B)   2.1 74.0 94.6 12.5 81.2 50.0   8.3 50.0 50.0   1.3
GPT-4o 89.2 83.3 91.6 10.5 77.0 51.9 72.8 53.7 54.7 55.1
AutoPresent 79.0 67.7 79.7 10.9 75.9 45.3   62.7 54.2 60.9 46.3
Code Generation w/ Expert-Designed Library
LLaVA (7B) 20.0 80.5 80.5   3.5 64.0 37.5 48.0 29.5 43.5   9.7
LLaMA (8B) 54.4 78.3 91.2 7.5 69.5 46.0 68.2 47.6 53.1 33.5
GPT-4o 86.7 86.2 92.5 12.7 76.3 54.6 83.7 70.5 59.4 58.0
AutoPresent (ours) 84.1 84.2 92.2 18.1 67.2 47.8 73.2 58.6 64.7 55.0
Table 9: Slide generation results (un-weighted by execution success) under the detailed instructions with images scenario.
E.2 Detailed Instructions Only
Table 10 shows all metrics down-weighted by the execution success rate; Table 11 shows reference-based and reference-free metrics without down-weighting by execution success.
Method ExecutionReference-Based Reference-Free Average
block text color position text image layout color
End-to-End Image Generation
Stable-Diffusion 100.0 74.5 33.4   9.0 75.0 19.6 45.1 36.9 40.5 48.0
DALLE 3 100.0 75.5 39.9   9.2 76.1 32.7 87.3 56.7 53.4 50.2
Code Generation w/o Library
LLaVA (7B) 17.9 12.2 16.3   1.4 12.4   7.9 15.3   5.7   5.0   9.5
LLaMA (8B) 4.6 63.0 87.0 17.4 80.4 30.4 19.6 41.3 47.8   2.8
GPT-4o 50.3 42.2 50.0   6.0 39.8 27.1 15.3 29.0 29.2 32.2
Code Generation w/ Expert-Designed Library
LLaVA (7B) 17.4 15.6 15.5   0.9 10.5   5.7   6.2   4.1   7.5   8.3
LLaMA (8B) 60.5 45.1 55.5   5.2 43.6 29.5 44.3 29.6 33.4 37.4
GPT-4o 87.7 72.3 80.8   6.0 65.9 46.6 73.0 58.5 52.9 56.3
AutoPresent (ours) 89.2 70.2 82.7   9.3 58.5 43.0 47.7 55.3 63.2 55.2
Table 10: Results (weighted by execution success) under detailed instructions only scenario.
Method ExecutionReference-Based Reference-Free Overall
block text color position text image layout color
End-to-End Image Generation
Stable-Diffusion 100.0 74.5 33.4   9.0 75.0 19.6 45.1 36.9 40.5 48.0
DALLE 3 100.0 75.5 39.9   9.2 76.1 32.7 87.3 56.7 53.4 50.2
Code Generation w/o Library
LLaVA (7B) 17.9 68.2 91.1   7.8 69.3 44.1 85.8 31.8 27.9 9.5
LLaMA (8B) 4.6 2.9   4.0   0.8   3.7   1.4   0.9   1.9   2.2   2.8
GPT-4o 50.3 83.9 92.4 11.9 79.1 53.9 30.4 57.7 58.1 32.2
Code Generation w/ Expert-Designed Library
LLaVA (7B) 17.4 89.7 89.1   5.2 60.3 32.8 35.6 23.6 43.1   8.3
LLaMA (8B) 60.5 74.5 91.7   8.6 72.1 48.8 73.2 29.6 48.9 37.4
GPT-4o 87.7 82.4 92.2   6.9 75.2 53.1 83.3 66.7 60.3 56.3
AutoPresent (ours) 89.2 78.7 92.7 10.4 65.6 48.2 53.5 62.0 70.9 55.2
Table 11: Results (un-weighted by execution success) under detailed instructions only scenario.
E.3 High-Level Instructions Challenge
Table 12 shows all metrics down-weighted by the execution success rate; Table 13 shows reference-based and reference-free metrics without down-weighting by execution success.
Method ExecutionReference-Based Reference-Free Average
block text color position text image layout color
End-to-End Image Generation
Stable-Diffusion 100.0 72.0 33.2   8.3 77.2   3.3 49.3 35.6 37.8 47.7
DALLE 3 100.0 73.5 48.2   7.6 77.3 14.9 89.7 57.2 52.4 51.7
CodeGen-based Methods w/o Library
LLaVA (7B) 19.5 14.9 13.2   1.7 13.6   8.0 16.8   5.9   6.2 10.0
LLaMA (8B) 8.7 7.6   6.3   0.7   4.7   4.6   2.4   5.0   5.4   4.8
GPT-4o 70.8 54.6 54.2   7.5 54.4 42.4 19.2 51.9 48.0 39.0
CodeGen-based Methods w/ Library
LLaVA (7B) 25.1 20.4 17.8   1.6 15.4   9.2   9.7   6.9 11.0 11.5
LLaMA (8B) 76.9 55.4 58.3   5.6 55.7 39.5 56.5 40.3 43.0 43.7
GPT-4o 97.4 77.0 75.8   7.7 73.7 59.7 73.8 78.7 65.4 58.5
AutoPresent (ours) 86.6 63.5 66.4 10.2 51.1 41.4 34.2 64.0 73.3 47.8
Table 12: Results (weighted by execution success) under high-level instructions scenario.
Method ExecutionReference-Based Reference-Free Average
block text color position text image layout color
End-to-End Image Generation
Stable-Diffusion 100.0 72.0 33.2   8.3 77.2   3.3 49.3 35.6 37.8 47.7
DALLE 3 100.0 73.5 48.2   7.6 77.3 14.9 89.7 57.2 52.4 51.7
CodeGen-based Methods w/o Library
LLaVA (7B) 19.5 76.4 67.7   8.7 69.7 41.0 86.2 30.3 31.8 10.0
LLaMA (8B)   8.7 87.4 72.4   8.0 54.0 52.9 27.6 57.5 62.1   4.8
GPT-4o 70.8 77.1 76.8 10.6 76.8 59.9 27.1 73.3 67.8 39.0
CodeGen-based Methods w/ Library
LLaVA (7B) 25.1 81.3 70.9   6.4 61.4 36.7 38.6 27.5 43.8 11.5
LLaMA (8B) 76.9 72.0 75.7   7.3 72.4 51.3 73.4 52.4 55.9 43.7
GPT-4o 97.4 79.0 77.8   7.9 75.6 61.3 75.8 80.7 67.1 58.5
AutoPresent (ours) 86.6 73.3 76.7 11.8 59.0 47.8 39.5 73.9 84.6 47.8
Table 13: Results (un-weighted by execution success) under high-level instructions scenario.
Appendix F Perceptual Analysis
In this section, we provide perceptual analysis details. We build a google doc and ask the user to score each slide from 1-5 (1 is the worst and 5 is the best), as shown in Figure 10.
Pleasescoreeachslidefrom1-5basedonyourpreferencetousethisslideinarealpresentation.5isthebest,1istheworst.
Carefullyreadingeachslide’scontentbeforeranking.
Figure 10: Instruction we used for the perceptual evaluation.
Figure 11: An example of the perceptual analysis question. We ask the human to score the quality of the slide from 1-5.
An example of the question is shown in Figure 11.